A student drops an object out the window of the top floor of a high-rise dormitory. (a) Neglecting air resistance, how fast is the object traveling when it strikes the ground at the end of 4.4 s? Express the speed in mi/h for a familiar comparison. mi/h Can things dropped from high places be dangerous to people below? (Assume that a falling object can be dangerous if its speed exceeds 35 mi/h.) Yes, an object dropped from a high place can accelerate to dangerous speeds. No, an object maintains a constant safe speed. Correct: Your answer is correct. (b) How far, in meters, does the object fall during the 4.4 s? m.
Answer #1

(A)

U = 0, t = 4.4 s

Let the speed after 4.4 s is v

Use first equation of motion

V = u + gt

V = 0 + 9.8 × 4.4

V = 43.12 m / s

Convert it into mph.

V = 96.46 mi/h

Yes it is dangerous.

(B)

Let it falls by distance h

Use second equation of motion

h = ut + 1/2 gt^2

h = 0 + 1/2 × 9.8 × 4.4 × 4.4

h = 94.864 m